# 1]AgiBot2]Shanghai AI Lab3]CUHK 4]SJTU 5]FDU 6]HKUST 7]HIT ]

### v2

@@ 3.3 From 4D Embodied Space to Physical Action
In addition to video generation, we integrate a policy head into the diffusion generator networks, enabling the generation of videos and corresponding actions after the extensive future space generation pretraining. The input latent vector of the policy head is extracted from the middle layer of the video diffusion generator, as shown in Fig. 1. To improve efficiency, this latent vector is taken from the noisiest step of the diffusion process, i.e., the feature vector after the first denoising step, reducing computational overhead. During policy prediction, the sparse memory stores visual images that are either observed or reconstructed FAVs under a multi-view setup as shown in Fig. 5. Actions are predicted in chunks, making the approach well-suited for time-sensitive robotic control tasks.
@@ 4 Experiments
4.1 Experiment Settings
Training Data. We selected several public datasets characterized by well-defined task logic, including RT-1 Brohan et al. (2022), Taco-Play Rosete-Beas et al. (2022), ManiSkill Gu et al. (2023), BridgeV1 Walke et al. (2023), LanguageTable Lynch et al. (2023) and RoboTurk Mandlekar et al. (2019), for pretraining. During pretraining , only video frames were utilized for video generation training. Furthermore, we constructed a dataset containing multi-anchor view video ground truths using the Isaac Sim simulator Mittal et al. (2023). The FAV generation model was trained by leveraging the weights derived from the single-view video generation model. For the policy planning task, fine-tuning with a limited quantity of demonstration data from specific scenarios proved sufficient to attain state-of-the-art performance. To mitigate domain gaps encountered when training with heterogeneous data, we employed domain embeddings inspired by Wang et al. (2024). Specifically, distinct domain embeddings were allocated to each sub-dataset. In subsequent space generation and policy planning, these embeddings were integrated with the diffusion timestep embeddings prior to input into the diffusion model. This methodology effectively alleviated conflicts arising from discrepancies in entities, task types, and visual styles.
Training Details. Our model is conducted based on UNet-based Video Diffusion Models (VDM) Xing et al. (2025), and can be easily adapted to DiT Peebles and Xie (2023b) architectures. In our experiments on generating embodied future spaces, we identified that chunk size significantly influences model performance. Comparative analyses utilizing chunk sizes of 1, 4, 8, and 16 revealed that the model exhibited optimal robustness when employing a chunk size of 8 (further details regarding these experiments can be found in the supplementary material). Following the methodology outlined in  Bruce et al. (2024), we introduced corruptive noise to the frames within the memory context. To alleviate degradation in autoregressive generation, the intensity of this noise was modulated in a cosine-related manner relative to the distance from the current moment. In the policy prediction experiment, the action head adopts the Diffusion Policy (DP) architecture Chi et al. (2023), with a total of 190M parameters. For the condition of the DP head, we utilize the feature before middle block of the UNet in the first denoise step, and calculate the mean value over spatial dimension, resulting in a final shape of $T\times C$, where $T$ is the length of video and $C$ is the number of channels before middle block. The rendered FAV images are with $512\times 320$ image sizes and the action header predicts the delta pose.
Figure 7: Qualitative results for multi anchor view generation on LIBERO (left) and real-world manipulation data (right), from AgiBot World AgiBot (2024). One view is overlapped with a fixed RGB sensor and other views are manully set. Visualized Frames are uniform. The consistency of objects across views by a red rectangle.
@@ 4.1 Experiment Settings
Diffusion Policy One Third View 78.3 92.5 68.3 50.5 72.4
Octo One Third View 78.9 85.7 84.6 51.1 75.1
OpenVLA One Third View 84.7 88.4 79.2 53.7 76.5
MDT One Third & One Wrist View 78.5 87.5 73.5 64.8 76.1
MAIL One Third & One Wrist View 74.3 90.1 81.8 78.6 83.5
EnerVerse One FAV 92.1 93.2 78.1 73.0 84.1
EnerVerse Three FAVs 91.2 97.7 85.0 80.0 88.5
@@
4.2 Comparison Results
Embodied Future Space Generation. Following AVID Rigter et al. (2024), we assess video generation quality utilizing the RT-1 Brohan et al. (2022) dataset. To create a comparable baseline, we fine-tune DynamicCrafter on the RT-1 dataset and run inference iteratively with FreeNoise Qiu et al. (2023) to enable long video generation(DC-FN). For evaluation, we generate 200 synthetic videos with varied lengths by conditioning the models on the initial frame and task instructions, subsequently comparing the generated videos against the ground truth using standard metrics such as PSNR and FVD. However, while these metrics primarily evaluate visual quality, embodied tasks necessitate additional considerations, including semantic alignment with instructions, workspace consistency across frames, and motion continuity. To address these higher-order aspects, we execute a user study involving robotics experts, assessing the generated videos based on semantic accuracy, frame consistency, and motion continuity.
Multi-View Consistency. In this section, we qualitatively demonstrate the capability of EnerVerse to generate multi-view videos of the same scene while ensuring consistency across anchor views. Furthermore, each view attains high-quality image generation, thereby highlighting the robustness of our approach. As shown in Fig. 7, EnerVerse could generate high-quality multi anchor view videos in both simulator and real-world settings.
Robotic Policy Evaluation Following the evaluation protocol in OpenVLA Kim et al. (2024), we evaluate robotic policies using the LIBERO Liu et al. (2024) benchmark, which consists of four distinct task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite contains 10 tasks, each with 50 human demonstrations. For each task suite, a separate policy model is fine-tuned. We compare our method against five baselines: Diffusion Policy Chi et al. (2023), a direct action learning policy trained from scratch; Octo Team et al. (2024), a transformer-based policy model fine-tuned on the target dataset; OpenVLA, a 7B vision-language-action (VLA) model fine-tuned on the target dataset; MDT Reuss et al. (2024),a diffusion transformer-based policy with an auxiliary MAE loss; MAIL Jia et al. (2024), a policy model with Mamba Gu and Dao (2023) in an encoder-decoder structure. For evaluation, all models are tested across tasks using 50 rollouts per task, with results averaged over three random seeds. Experiments with EnerVerse-A are conducted under two setups: a single static-camera, consistent with OpenVLA, and a three-static-camera configuration as shown in Fig. 5.
As shown in Tab. 2, EnerVerse achieves state-of-the-art performance across the LIBERO benchmark, significantly surpassing all baselines. With a One Third View input, it achieves an average score of 84.0, outperforming strong baselines like MAIL (83.5) and OpenVLA (76.5). The Three Third View configuration further enhances performance, achieving the highest average score of 88.5, demonstrating the value of richer visual input. The model’s balanced performance across all tasks, particularly excelling in Object and Goal tasks, underscores its robustness and adaptability.
@@ 4.3 Further Studies
Sparse Memory Mechanism. We evaluate the effectiveness of our sparse memory mechanism in both policy learning and video generation. The evaluation is conducted on the LIBERO-Long task suite, as this suite involves significantly longer task execution steps, requiring the policy to exhibit strong long-range memory and task reasoning capabilities. The evaluation is performed with a single visual input. As shown in Tab. 3, the absence of the sparse memory results in significant performance degradation, with the policy achieving only 30.8 compared to 73 when the sparse memory mechanism is applied. Similarly, Fig. 8 demonstrates that when the video generator operates without sparse memory, the model experiences unexpected collapse and fails to recover in out-of-distribution (OOD) scenarios. In contrast, the sparse memory mechanism ensures robust performance while also saving computational resources.
@@
Training Strategy Analysis. To analyze the impact of different training strategies on robotic policy learning, we trained four robotic policies on the LIBERO-Spatial task suite using the following approaches: (1) training the entire EnerVerse from scratch using only policy loss optimization; (2) training the entire EnerVerse as in (1) but initialized with pretrained weights from a general video generator, e.g. DynamiCrafter(DC) Xing et al. (2025), which is trained with the general natural videos; (3) co-training EnerVerse by optimizing both the robotic policy action loss and the video generation loss simultaneously; and (4) the default two-stage training strategy, where the video generator is pretrained first, followed by fine-tuning EnerVerse using only robotic policy loss optimization.
@@
Attention Map Analysis. To further analyze the alignment between the predicted action space and the visual space, including the visual observations cached by our Sparse Memory Mechanism and the generated future space, we visualized the attention maps from the first several layers of the Cross-Attention Block in our policy head.
Real-World Experiments. To evaluate the manipulation capabilities of EnerVerse, we conducted real-world experiments using commercial robotics in two challenging industrial scenarios. Unlike the evaluation tasks described in Sec. 4.2, these scenarios required precise manipulation and robust decision-making. In the first task, the robot placed blocks into designated compartments of a foam worktable, demanding accuracy due to the tight fit and visual similarity between the foam and table. In the second task, the robot sorted several transparent plastic objects, including a measuring cup and plate, where the transparency added complexity to object recognition and manipulation. For details, please refer to Appendix F.
5 Conclusion
@@
In conclusion, EnerVerse is a generative robotics foundation model that tackles multi-view video generation and long-range policy execution by modeling embodied future spaces. With sparse contextual memory and Free Anchor Views (FAVs), EnerVerse enhances spatial reasoning and task adaptability. The EnerVerse-D pipeline, combining generative modeling with 4DGS, bridges the sim-to-real gap, reducing reliance on real-world data. Integrated with a policy head, EnerVerse-A achieves state-of-the-art performance in manipulation tasks.
F Real-World Robotic Experiments

### v3

@@ 2.3 From 4D Embodied Space to Physical Action
We further integrate a policy head into the diffusion generator, enabling the generation of actions after the extensive pretraining of future space generation. The policy operates on a compact visual latent $E$ extracted from the middle block of the UNet backbone at the first denoising step (the noisiest step), which reduces computation while retaining rich task-relevant cues. Visual inputs may be captured RGB frames or rendered views, as shown in Fig 4; both are encoded by the shared video backbone. The action head $h_{\theta}$ is a stack of DiT blocks followed by a linear projection to the action space. We predict action chunks zhao2023learning for temporal consistency. Let $a_{t:t+\tau-1}\in\mathbb{R}^{\tau\times d}$ denote a chunk of actions with $d=7$ (delta position, rotation, and gripper openness). The denoising model $f_{\theta}$ estimates clean actions from noisy inputs using DDPM-style training:
$a_{t:t+\tau-1}^{\,0}\leftarrow f_{\theta}(c,o_{t},a_{t:t+\tau-1}^{\,k},k)\;=\;h_{\theta}(E,a_{t:t+\tau-1}^{\,k},k),$
where $E$ is the cached visual latent from EnerVerse-G and $k\in\{1,\ldots,K\}$ is the diffusion step. The training objective minimizes denoising MSE.
At inference, we compute $E$ once by passing $(c,o_{t})$ through the video diffusion backbone and cache it across the action denoising steps. The action head then iteratively denoises from $a^{K}$ to $a^{0}$ for the current chunk. We use per-view, per-frame decoding for visuals, but policy conditioning is view-aggregated via $E$ (mean over spatial dimensions within the UNet middle block). We provide more details in the Appendix.
Figure 4: Render View 1 and Render View 2 are generated by rendering from a point cloud reconstructed from RGB-Image 1 using depth wrapping. The render views correspond to camera views obtained by rotating the RGB camera view around the Z-axis by $\pm 30^{\circ}$.
@@ 3 Experiments
3.1 Experiment Settings
Training Data: We selected several public datasets characterized by well-defined task logic, including RT-1 brohan2022rt , Taco-Play rosete2022tacorl , ManiSkill gu2023maniskill2 , BridgeV1 walke2023bridgedata , LanguageTable lynch2023interactive , and RoboTurk mandlekar2019scaling for pretraining. Furthermore, we constructed a dataset containing multi-view video ground truths using the Isaac Sim simulator mittal2023orbit . The detailed dataset statistics could be found in Appendix. During pretraining , only video frames were utilized for video generation training. For the policy planning task, fine-tuning with a limited quantity of demonstration data from specific scenarios proved sufficient to attain state-of-the-art performance. To mitigate domain gaps encountered when training with heterogeneous data, we employed domain embeddings inspired by wang2024scaling . Specifically, distinct domain embeddings were allocated to each sub-dataset. In subsequent space generation and policy planning, these embeddings were integrated with the diffusion timestep embeddings prior to input into the diffusion model. This methodology effectively alleviated conflicts arising from discrepancies in entities, task types, and visual styles.
Model Details. Our model is conducted based on UNet-based Video Diffusion Models (VDM) xing2025dynamicrafter , and can be easily adapted to DiT peebles2023scalablediffusionmodelstransformers architectures. And the image decoding occurs per-view and per-frame. In our experiments on generating embodied future spaces, we identified that chunk size significantly influences model performance. Comparative analyses utilizing chunk sizes of 1, 4, 8, and 16 revealed that the model exhibited optimal robustness when employing a chunk size of 8. Following the methodology outlined in  bruce2024genie , we introduced corruptive noise to the frames within the memory context. To alleviate degradation in autoregressive generation, the intensity of this noise was modulated in a cosine-related manner relative to the distance from the current moment. After pretraining the multiview video generation models, we performed a generation learning with the action videos to achieve both visual and spatial adaptation. Subsequently, we fine-tuned the action policy head using action trajectories. Following this, we fine-tune the action policy head using the action trajectories. The action head adopts the Diffusion Policy (DP) architecture chi2023diffusion , with a total of 190M parameters. For the DP head’s condition, we utilized features extracted from the middle block of the UNet during the first denoising step and calculated the mean across the spatial dimensions. This resulted in a final shape of $T\times C$, where $T$ is the video length and $C$ is the channel number. Rendered images have a resolution of $512\times 320$, and the action head predicts the delta pose. For EnerVerse-D, we integrate 4D Gaussian Splatting using the official implementation wu20244d , with depth-based initialization when available and a deformation depth of 1.
3.2 Comparison Results
Embodied Future Space Generation. Following AVID rigter2024avid , we assess video generation quality utilizing the RT-1 brohan2022rt dataset. To create a comparable baseline, we fine-tune DynamicCrafter on the RT-1 dataset and run inference iteratively with FreeNoise qiu2023freenoise to enable long video generation(DC-FN). For evaluation, we generate 200 synthetic videos with varied lengths by conditioning the models on the initial frame and task instructions, subsequently comparing the generated videos against the ground truth using standard metrics such as PSNR and FVD. However, while these metrics primarily evaluate visual quality, embodied tasks necessitate additional considerations, including semantic alignment with instructions, workspace consistency across frames, and motion continuity. To address these higher-order aspects, we execute a user study involving robotics experts, assessing the generated videos based on semantic accuracy, frame consistency, and motion continuity.
@@ 3.2 Comparison Results
Multi-View Generation Consistency. In this section, we qualitatively demonstrate the capability of EnerVerse to generate multi-view videos of the same scene while ensuring consistency across views. Furthermore, each view attains high-quality image generation, thereby highlighting the robustness of our approach. As shown in Fig. 6, EnerVerse could generate high-quality multi-view videos in both simulator and real-world settings.
Figure 6: Qualitative results for multi view generation on LIBERO (left) and real-world manipulation data (right). The first generated view position is overlapped with a static mounted RGB camera and others are manually set. The consistency of objects across views is highlighted by a red rectangle.
Robotic Policy Evaluation on LIBERO Following the evaluation protocol in OpenVLA kim2024openvla , we evaluate robotic policies using the LIBERO liu2024libero benchmark, which consists of four distinct task suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite contains 10 tasks, each with 50 human demonstrations. For each task suite, a separate policy model is fine-tuned. We compare our method against five baselines: Diffusion Policy chi2023diffusion , a direct action learning policy trained from scratch; Octo team2024octo , a transformer-based policy model fine-tuned on the target dataset; OpenVLA, a 7B vision-language-action (VLA) model fine-tuned on the target dataset; MDT reuss2024multimodal ,a diffusion transformer-based policy with an auxiliary MAE loss; MAIL jia2024mail , a policy model with Mamba gu2023mamba in an encoder-decoder structure. Besides, we provide the results of MAIL with two S-RGB input with their official implementation. For evaluation, all models are tested across tasks using 50 rollouts per task, with results averaged over three random seeds. Experiments with EnerVerse-A are conducted under three setups: a single RGB image, and when RGB-D is available, 1 RGB image with 1 rendered view, and 1 RGB image with 2 rendered views, as shown in Fig. 4. The abbreviations denote different input modalities: S-RGB for Static RGB, G-RGB for Gripper RGB, S-RGBD for Static RGB-D, G-RGBD for Gripper RGB-D, P for proprioceptive arm position.
@@
Diffusion Policy S-RGB 78.3 92.5 68.3 50.5 72.4
Octo S-RGB 78.9 85.7 84.6 51.1 75.1
OpenVLA S-RGB 84.7 88.4 79.2 53.7 76.5
MDT S-RGB,G-RGB 78.5 87.5 73.5 64.8 76.1
MAIL S-RGB,G-RGB 74.3 90.1 81.8 78.6 81.2
MAIL S-RGB,S-RGB 76.0 90.0 82.0 78.0 81.5
EnerVerse S-RGB 92.1 93.2 78.1 73.0 84.1
EnerVerse S-RGBD $\rightarrow$ RGB with 1 Render 93 95.0 81.0 73.0 85.5
EnerVerse S-RGBD $\rightarrow$ RGB with 2 Render 91.2 97.7 85.0 80.0 88.5
@@
As shown in Tab. 2, EnerVerse achieves state-of-the-art performance across the LIBERO benchmark, significantly surpassing all baselines. With single S-RGB input, it achieves an average score of 84.1, outperforming strong baselines.
[…]
Method Input 1 2 3 4 5 Avg. Len.
RoboFlamingo li2023vision S-RGB, G-RGB 82.4 61.9 46.6 33.1 23.5 2.47
GR-1 wu2023unleashing S-RGB, G-RGB, P 85.4 71.2 59.6 49.7 40.1 3.06
3D Diffuser ke20243d S-RGBD, G-RGBD, P 92.2 78.7 63.9 51.2 41.2 3.27
SUSIE black2023zero S-RGB 87 69.0 49.0 38.0 26.0 2.69
EnerVerse S-RGB 90.8 73.0 57.3 43.7 35.6 3.00
Table 3: The comparisons with state-of-the-art approaches on Calvin (ABC $\rightarrow$ D).
@@ 3.3 Further Studies
Effectiveness of Sparse Memory Mechanism. We evaluate the effectiveness of our sparse memory mechanism in both policy learning and video generation. The evaluation is conducted on the LIBERO-Long task suite, as this suite involves significantly longer task execution steps, requiring the policy to exhibit strong long-range memory and task reasoning capabilities. The evaluation is performed with a single visual input. As shown in Tab. 4, the absence of the sparse memory results in significant performance degradation, with the policy achieving only 30.8 compared to 73 when the sparse memory mechanism is applied. Similarly, Fig. 7 demonstrates that when the video generator operates without sparse memory, the model experiences unexpected collapse and fails to recover in out-of-distribution (OOD) scenarios. In contrast, the sparse memory mechanism ensures robust performance while also saving computational resources.
@@
Training Strategy Analysis. To analyze the impact of different training strategies on robotic policy learning, we trained four robotic policies on the LIBERO-Spatial task suite using the following approaches: (1) training the entire EnerVerse from scratch using only policy loss optimization; (2) training the entire EnerVerse as in (1) but initialized with pretrained weights from a general video generator, e.g. DynamiCrafter(DC) xing2025dynamicrafter , which is trained with the general natural videos; (3) co-training EnerVerse by optimizing both the robotic policy action loss and the video generation loss simultaneously; and (4) the default two-stage training strategy, where the video generator is pretrained first, followed by fine-tuning EnerVerse using only robotic policy loss optimization.
@@
Effectiveness of Additional Rendered Views. With the expressive pretrained generative space prior, our method using a single S-RGB camera already achieves SOTA performance. When a single S-RGBD camera is available, we can incorporate additional rendered views, with the original RGB image as input to the model. These additional views not only bring the input setting closer to the training configuration but also enable the policy model to better leverage the pretrained generative space prior. Notably, the configuration RGB with 1 Render outperforms MAIL’s 2 S-RGB setup, both in terms of overall performance and gains compared to a single RGB input, demonstrating that the performance improvement is not solely due to the additional visual inputs. Incorporating RGB with 2 Render yields even greater gains by mitigating occlusions, as illustrated in Fig. 4.
3D Video vs. 4D Space for Robotics. We provide a direct comparison between attaching a diffusion policy head to a base single-view video generator (DynamiCrafter) and our 4D variant (EnerVerse-A). As shown in Table 6, the base video model underperforms substantially (79.0) relative to EnerVerse-A (92.1), highlighting the benefits of 4D extensions. We hypothesize that the cross-view consistency learned during multi-view pretraining provides stronger geometric priors, which help the model better understand spatial relationships and occlusions. Further, even when tested with single RGB-D inputs, EnerVerse benefits from additional rendered views at inference. These findings underscore the ability to incorporate additional rendered views at test time further enhances performance, showcasing the practicality and effectiveness of our approach.
Model Multi-view Pre-Train Input at Test SR
DynamiCrafter + DP No S-RGB 79.0
EnerVerse-A Yes S-RGB 92.1
EnerVerse-A Yes S-RGB with 1 Render 93.0
Table 6: Direct comparison between a base single-view video generator with a policy head and our approach on LIBERO-Spatial. Multi-view pretraining substantially improves SR even with single-view inputs.
Real-World Experiments. To evaluate the manipulation capabilities of EnerVerse-A, we conducted real-world experiments using commercial robotics on the tasks of Block Placing, Plastic Objects Sorting, and Fruit Sorting. For further details, please refer to Appendices A.
More Discussions and Experiments. We provide additional discussions and experiments in the appendices: pretrained model performance (Appendices C), alignment between action and visual spaces (Appendices D), robustness against OOD samples (Appendices E), model architecture and computational overhead (Appendices F and G), and visual samples validating the data engine’s effectiveness (Appendices I).
4 Related Works
Video Generation Models. Diffusion-based video generation models have made notable progress, especially in text-to-video (T2V) generation blattmann2023stable ; song2020denoising . Early T2V approaches 2023i2vgenxl ; chen2023videocrafter1 ; ren2024consisti2v ; guo2023animatediff build on text-to-image (T2I) priors by introducing temporal modules trained on video data. DynamicCrafter xing2025dynamicrafter reuses motion priors from T2V diffusion models in an image-to-video (I2V) context. Recent works  kong2024hunyuanvideo ; opensora ; bao2024vidu explores replacing U-Nets with Diffusion Transformer (DiT) peebles2023scalable , and  ren2025gen3c uses the multi-camera poses information to extend the video generation into the 3D world modelling. Other studies gao2024vid incorporate causal mechanisms to generate longer sequences or extend video-generation models into world modeling by forecasting future states hu2023gaia ; bruce2024genie ; wang2023drivedreamer . In this paper, we mainly adopt DynamicCrafter as our base I2V framework due to its open-source availability and widespread use. We also ensure compatibility with modern DiT architectures, although that is not our main focus here.
Video Pretraining for Robotics. GR-2 cheang2024gr presents a generalizable robot manipulation framework that pretrains on large-scale internet videos, then fine-tunes on both video generation and action prediction for robotic trajectories. LAPA ye2024latent uses non-robot action videos for representation learning, mapping discrete latent actions (via VQ-VAE) to robotic manipulation tasks through a VLA model. SEER tian2024seer further explores inverse dynamics pretraining to boost performance. AVID rigter2024avid employs DynamicCrafter xing2025dynamicrafter as its foundation, using an adapter for the robotics domain. VidMan wen2024vidman , based on OpenSora opensora , focuses on environment prediction before action generation but is limited to 2D image space. In contrast, we propose generating long-sequence futures via a novel data generation engine, capturing richer motion information vital for robotics.
4D Generation. Recent progress chen2024survey allows reconstruction of dynamic scenes from 2D videos using 3D Gaussian Splatting (GS) kerbl20233d and Neural Radiance Fields (NeRF) mildenhall2021nerf . Prior approaches approximate the spatio-temporal 4D volume with sets of 4D Gaussians yang2023real , jointly optimizing geometry and motion in canonical space wu20244d . More recent advancements li2024vivid employ customized sampling for multi-view video diffusion models, particularly for single dynamic objects. DimensionX sun2024dimensionx leverages multiple LoRAs hu2021lora for diverse camera motions, while Cat4D wu2024cat4d uses a single multi-view diffusion model to generate videos for dynamic 3D reconstruction. By contrast, our method produces videos from a multi-views tailored for robotic manipulation tasks. In our offline data flywheel stage, GS complements video generation models to mitigate the sim-to-real gap, enhancing geometric consistency and reducing hallucinations from the generative models.
5 Conclusions
@@
In conclusion, EnerVerse is a generative robotics foundation model that tackles multi-view video generation and long-range policy execution by modeling embodied future spaces. With sparse contextual memory and chunkwise autoregressive architecture, EnerVerse enhances spatial reasoning and task adaptability. The EnerVerse-D pipeline, combining generative modeling with 4DGS, bridges the sim-to-real gap, reducing reliance on real-world data. Integrated with a policy head, EnerVerse-A achieves state-of-the-art performance in manipulation tasks, both in the simulator environment and the real-world settings.
Limitations. Due to the high dynamics and rich object interactions in robotics tasks, video generation models inevitably produce artifacts, as discussed in App. B. However, advancements in the video modeling community are expected to improve this. Additionally, while we provide an initial attention map analysis, further exploration is needed to better understand how video generation guides action policy learning. Finally, the current rendered views are derived from heuristically set camera poses, which may not be optimal. Integrating Next-Best View methods zhang2023affordance could address this limitation.
Appendix A Real-World Robotic Experiments